A family owns two corner lots of land that form a trapezoid with bases of 200 feet and 175 feet and a height of 60 feet. They decide to sell the triangular piece of land. How much land will they have left after the sale?

Answer by addingup(3677)

area of the trapezoid:
(a + b)/2 x h
= (200 + 175)/2 x 60
= 187.50 x 60
= 11,250 this is the area of the trapezoid
----------------------------------
area of the triangle:
bh/2
the base is 200 - 175 = 25
(25 x 60)/2
= 1500 this is the area of the triangle.
:
Subtract the area of the triangle from the area of the trapezoid to get your answer.